# Conditional Generative Representation for Black-Box Optimization with Implicit Constraints

### v3

@@ Abstract
1 Introduction
In recent years, black-box optimization (BBO) has emerged as a critical approach in addressing complex decision-making challenges across various domains, particularly when dealing with objective functions that are difficult to analyze or explicitly define. Unlike traditional optimization methods that require gradient information or explicit mathematical formulations, black-box optimization treats the objective function as a “black box” that can be queried for function values but offers no additional information about its structure [32]. This methodology is especially valuable for designing public policy, such as police redistricting [52, 53], site selection for emergency service systems [48], hazard assessment [47] and public healthcare policymaking [5]. Policymakers and researchers in these areas frequently encounter optimization problems embedded within complex human systems, where decision evaluations are inherently implicit, and conducting them can be resource-intensive. Black-box optimization, therefore, stands out as a potent tool for navigating through these intricate decision spaces, providing a means to optimize outcomes without necessitating a detailed understanding of the underlying objective function’s analytical properties.
@@ 1 Introduction
However, the broader application prospects of BBO in public policymaking are hampered due to two major hurdles: (1) Defining the feasible region or setting clear constraints for decisions is inherently complicated for real human systems. Policymakers often encounter a myriad of both explicit and implicit rules when making an optimal decision, adding a significant layer of complexity to the optimization process. For instance, in police districting, police departments often organize their patrol forces by dividing the geographical region of a city into multiple patrol areas called zones. Their goal is to search for the optimal districting plan that minimizes the workload variance across zones [25, 52, 53]. As shown in Figure 1, a well-conceived plan necessitates each zone to adhere to certain shape constraints (e.g., contiguity and compactness) that are analytically challenging to formulate [36], while also taking socio-economic or political considerations into account (e.g., ensuring fair access to public facilities). This creates a web of implicit constraints [6, 3] that are elusive to define clearly and have a complex high-dimensional structure (e.g., manifold shape), making the assessment of feasible region nearly as expensive as evaluating the objective itself [11]. (2) The decisions for public policy are usually high-dimensional, presenting a significant computational hurdle to utilizing traditional BBO methods [29, 45, 4]. For example, the redistricting problem can be formulated by mixed-integer programming, grappling with hundreds or even thousands of decision variables even for medium-sized service systems [52, 53].
@@
Despite the difficulty in formulating constraints for decisions in public policymaking, the rich repository of historical decisions adopted by the practitioners, combined with the increasingly easier access to human systems [43, 49], offer a wealth of decision samples. Collecting these samples might involve seeking guidance from official public entities on decision feasibility or generating decisions grounded in domain expertise, bypassing the need to understand the explicit form of the constraints. These readily available decisions harbor implicit knowledge that adeptly captures the dynamics of implicit constraints, providing a unique opportunity to skillfully address these issues. This inspires us to develop a conditional generative representation model that maps the feasible region in the original space to a lower-dimensional latent space, which encapsulates the key pattern of these implicit constraints. As a result, the majority of existing BBO methods can be directly applied to solve the original optimization problem in this latent space without constraints.
@@
In this paper, we aim to solve Implicit-Constrained Black-Box Optimization (ICBBO) problems. In these problems, the constraints are not analytically defined, however the feasibility of a given decision can be easily verified. We introduce a new approach called Conditional And Generative Black-box Optimization (CageBO). Using a set of labeled decisions as feasible or infeasible, we first construct a conditional generative representation model based on the conditional variational autoencoder (CVAE) [38] to explore and generate new feasible decisions within the intricate feasible region. To mitigate the impact of potentially poorly generated decisions, we adopt a post-decoding process that aligns these decisions closer to the nearest feasible ones. Furthermore, we incorporate this conditional generative representation model as an additional surrogate into our black-box optimization algorithm, allowing for a two-way mapping between the feasible region in the original space and an unconstrained latent space. As a result, while objective function is evaluated in the original space, the black box optimization algorithm is performed in the latent space without constraints. We prove that our proposed algorithm can achieve a no-regret upper bound with a judiciously chosen number of observations in the original space. Finally, we validate our approach with numerical experiments on both synthetic and real-world datasets, including applying our method to address large-scale redistricting challenges within police operation systems in Atlanta, Georgia, demonstrating its significant empirical performance and efficiency over existing methods.
@@
    We formulate a novel class of optimization problems called Implicit Constrained Black-Box Optimization (ICBBO) and develop the CageBO algorithm that effectively tackles high-dimensional ICBBO problems.
@@
    We introduce a conditional generative representation model that constructs a lower-dimensional, constraint-free latent space. This space enables BBO algorithms to efficiently search and generate new feasible candidate solutions.
@@
    By proving the no-regret expected cumulative regret bound, we show that our CageBO algorithm is able to find the global optimal solution.
@@
    Our empirical results demonstrate the superior performance of our model against the baseline methods on both synthetic and real data sets, particularly in scenarios with complex implicit constraints.
@@
Related work. Black-box optimization, a.k.a. zeroth-order optimization or derivative-free optimization, is a long-standing challenging problem in optimization and machine learning. Existing work either assumes the underlying objective function is drawn from some Gaussian process [46] or some parametric function class [7, 28]. The former one is usually known as Bayesian optimization (BO), with the Gaussian process serving as the predominant surrogate model. BO has been widely used in many applications, including but not limited to neural network hyperparameter tuning [19, 41], material design [42, 51], chemical reactions [16], and public policy [48].
@@
In numerous real-world problems, optimization is subject to various types of constraints. Eriksson proposes scalable BO with known constraints in high dimensions [10]. Letham explores BO in experiments featuring noisy constraints [26]. Gelbart pioneered the concept of BO with unknown constraints [12], later enhanced by Aria through the ADMM framework [2]. Their constraints are unknown due to uncertainty but can be evaluated using probabilistic models. In addition, Choi and Audet study the unrelaxable hidden constraints in a similar way [6, 3], where the feasibility of a decision can be evaluated by another black-box function. In contrast, the implicit constraints are unknown due to the lack of analytical formulations in our problem.
@@
Building on latent space methodologies, Varol presented a constrained latent variable model integrating prior knowledge [44]. Eissman [9] presents a VAE-guided Bayesian optimization algorithm with attribute adjustment. Deshwal and Doppa focus on combining latent space and structured kernels over combinatorial spaces [8]. Maus further investigates structured inputs in local spaces [30], and Antonova introduces dynamic compression within variational contexts [1]. However, it’s worth noting that none of these studies consider any types of constraints in their methodologies.
Additionally, there is another line of work aiming to solve black-box optimization in the context of chemical design. Notable contributions in this domain include [13, 50, 15]. They focus on the efficient exploration and optimization of new molecules using different generative models. Most recently, an active line of research has been exploring the use of diffusion models to address black-box optimization via sampling techniques, as detailed in pioneering studies [24, 23, 27].
2 Preliminaries
@@
Bayesian optimization. The BO algorithms prove especially valuable in scenarios where the evaluation of the objective function is costly or time-consuming, or when the gradient is unavailable. This approach revolves around constructing a surrogate model of the objective function and subsequently employing an acquisition function based on this surrogate model to determine the next solution for evaluation. For the minimization problem, a popular choice of surrogate model is the Gaussian process with the lower confidence bound (LCB) [40] serving as the acquisition function.
@@ 3 Proposed method
3.1 Conditional generative representation
@@
Suppose there exists a joint distribution between the decision variable $x\in\mathcal{X}$, which can be defined in a high-dimensional hybrid space of discrete and continuous variables, and a continuous latent variable $z\in\mathcal{Z}\subseteq\mathbb{R}^{d}$, conditioned on the feasibility criterion $c\in\{0,1\}$. We model the conditional distributions $q_{\phi}(z|x,c)$, $p_{\theta}(x|z,c)$, and the conditional prior $p(z|c)$ through the use of neural networks as outlined in [33]. Here, $\phi$ and $\theta$ represent the neural network weights associated with each respective distribution. Since it is intractable to directly optimize the marginal maximum likelihood of $x$, the evidence lower bound (ELBO) [18] of the log-likelihood is derived as follows
$\begin{split}\mathcal{L}_{\mathrm{ELBO}}=w(c)\underset{q_{\phi}(z|x,c)}{\mathbb{E}}\left[\log p_{\theta}(x|z,c)\right]-\eta\mathrm{D}_{\mathrm{KL}}\left(q_{\phi}(z|x,c)\|p(z|c)\right),\end{split}$ (1)
@@
The post-decoding process is initiated only for decoder-generated decisions $x$ that are infeasible. Feasible decisions are directly incorporated into the observed feasible set $\mathcal{D}_{f}$. Through iteratively expanding this observed feasible set, we enhance the accuracy of this process by improving the coverage of $\mathcal{D}_{f}$ on the underlying feasible region $\mathcal{S}$. This is further detailed in the ablation study in Appendix C.
3.2 Surrogate model
@@
We use a GP as our surrogate model of the indirect objective function $g$, denoted by $\text{GP}({\mu},{k};\mathcal{Z})$. In our problem, the mean function can be written as ${\mu(z)=\mathbb{E}[g(z)]}$. In addition, we adopt the Matérn kernel [34] as the kernel function $k(z,z^{\prime})=\mathbb{E}[(g(z)-\mu(z))(g(z^{\prime})-\mu(z^{\prime}))]$, which is widely-used in BO literature. The main advantage of the GP as a surrogate model is that it can produce estimates of the mean evaluation and variance of a new latent variable, which can be used to model uncertainty and confidence levels for the acquisition function described in the following. Note that the latent variable $z$ is assumed to follow a Gaussian prior in the latent decision model, which aligns with the assumption of the GP model that the observed latent variables $Z$ follow the multivariate Gaussian distribution.
@@
Figure 3: Solution paths of the same optimization problem suggested by our CageBO algorithm in the (a) original and (b) latent spaces, respectively. The objective in this illustrative example is to minimize the Ackley function subject to an “implicit” constraint (a circle) in a 2-dimensional decision space.
@@ 3.3 Theoretical analysis
We assume the distance between any two points in $\mathcal{Z}$ can be upper bounded by their distance in $\mathcal{X}$, i.e., $\forall x,x^{\prime}\in\mathcal{X},\|\mathtt{enc}(x)-\mathtt{enc}(x^{\prime})\|_{2}\leq C_{p}\|x-x^{\prime}\|_{2}$. We further assume that function $g:\mathcal{Z}\mapsto\mathbbm{R}$ is drawn from a Gaussian Process prior and it is $C_{g}$-Lipschitz continuous, i.e., $\forall z,z^{\prime}\in\mathcal{Z},|g(z)-g(z^{\prime})|\leq C_{g}\|z-z^{\prime}\|_{2}$. Here $C_{p}$ and $C_{g}$ are universal constants. We use big $\tilde{O}$ notation to omit any constant and logarithmic terms. Now we are ready to state our main theoretical result.
@@
We present an upper bound of expected cumulative regret of our Algorithm 1. This is a no-regret algorithm since $\lim_{T\rightarrow\infty}\mathbb{E}[R_{T}]/T=0$ which means our algorithm is able to find the global optimal solution by expectation. The bound has two terms. The first term follows from GP-UCB [39] where the maximum information gain depends on the choice of kernel used in the algorithm. For linear kernel, $\gamma_{T}=O(d\log T)$ and for squared exponential kernel, $\gamma_{T}=O(\log T)^{d+1}$. The second term is the regret term incurred by the post decoder $\psi_{\mathcal{D}_{f}}$ when needed, which is also sublinear in $T$ if $n$ is chosen no larger than $T$. Due to page limit, full proof of Theorem 1 is shown in Appendix A.
4 Experiments
4.1 Experimental setup
We compare our CageBO algorithm against three baseline approaches that can be used to address ICBBO problems. These baselines include (1) vanilla version of Bayesian optimization (BO) [17], (2) simulated annealing (SA) [22], (3) approximated Mixed Integer Linear Programming (MILP) [53], (4) a VAE-guided Bayesian optimization (VAE-BO) [9], and (5) a variant of constrained VAE-guided Bayesian optimization (CON-BO) [14]. We train the latent decision model in our framework with the Adam optimizer [21]. Both CageBO, VAE-BO, and CON-BO are trained and performed under an identical environment. MILP is only compared in the redistricting problems because their implicit constraints can be approximated by a set of linear constraints, albeit with the trade-offs of adding auxiliary variables and incurring computational expenses. Each method is executed 10 times across all experiments to determine the $95\%$ confidence interval of their results.
@@ 4.1 Experimental setup
4.2 Synthetic results
(a) Keane’s bump func.
(b) Michalewicz func.
@@ 4.2 Synthetic results
We consider minimizing (a) the $30$-d Keane’s bump function [20] and (b) the $30$-d Michalewicz function [31], both of which are common test functions for constrained optimization. To obtain samples from implicit constraints, the conundrum we aim to address via our methodology, we first generate $n=2,000$ samples from the standard uniform distribution in the $10$-dimensional latent space, then decode half of them through a randomly initialized decoder and mark those as feasible. We define the feasibility oracle to return 1 only if the input solution is matched with a feasible solution. We scale the samples accordingly so that the test functions can be evaluated under their standard domains, i.e., the Keane’s bump function on $[0,10]^{d}$ and the Michalewicz function on $[0,\pi]^{d}$.
@@
4.3 Case study: Police redistricting
@@
In this equation, ${\tau}_{j}$ represents the average travel time within district $j$, which can be estimated using the hypercube queuing model [25], which can be regarded as a costly black-box function. Assuming $\mu$ represents the uniform service rate across all districts and $\lambda_{j}$ denotes the arrival rate in district $j$. Essentially, $\rho_{j}$ quantifies the cumulative working duration per unit time for police units in district $j$. For clarity, a value of $\rho_{j}=10$ implies that the combined working time of all police units in district $j$ counts to 10 hours or minutes in every given hour or minute.
(a) MILP: $4.16(\pm 0.01)$
(b) SA: $2.96(\pm 0.66)$
(c) VAE-BO: $1.51(\pm 0.38)$
(d) CageBO: $1.42(\pm 0.31)$
(e) MILP: $3.27(\pm 0.01)$
(f) SA: $3.19(\pm 0.35)$
(g) VAE-BO: $2.32(\pm 0.06)$
(h) CageBO: $2.17(\pm 0.12)$
Figure 5: Districting plans with the corresponding workload variance (with 95% CI) for the police redistricting problem in a synthetic $6\times 6$ grid and a real-world scenario in Atlanta, Georgia. The workload in each region is indicated by the depth of the color. The CageBO plan achieves the lowest workload variance among all districts, indicating an optimal distribution of workload.
@@
Atlanta police redistricting. For the police zone design in Atlanta, there are 78 police service regions that need to be divided into 6 zones, with the decision variable represented as $x\in\{0,1\}^{78\times 6}$. The redistricting of Atlanta’s police zones is constrained by several potential implicit factors. These include contiguity and compactness, as well as other practical constraints that cannot be explicitly defined. One such implicit constraint is the need for changes to the existing plan should be taken in certain local areas. A drastic design change is undesirable because: (1) A large-scale operational change will result in high implementation costs. (2) A radical design change will usually face significant uncertainties and unpredictable risks in future operations. The arrival rate $\lambda_{l}$ and the service rate $\mu$ are estimated using historical 911-calls-for-service data collected in the years 2021 and 2022 [54, 55, 53]. We generate an initial data set $\mathcal{D}$ by simulating $n=10,000$ labeled decisions which are the neighbors of the existing plan with changes in certain local areas.
@@
Figure 4(d) displays the convergence of our algorithm in comparison to other baseline methods, showing a consistent reduction in workload variance. Notably, the redistricting plan created by our CageBO algorithm most closely aligns with the actual heuristic policy applied by policymakers, wherein the zone with the highest arrival rate is assigned the largest workload, and the workloads of other zones are well-balanced. As shown in Figure 5, the plan produced by our CageBO algorithm achieves the lowest workload variance, surpassing the baseline algorithms. The superior performance of our CageBO algorithm highlights its capacity to offer critical managerial insights to policymakers, thereby facilitating informed decision-making in real-world applications.
5 Conclusion

### v4

# Black-Box Optimization with Implicit Constraints for Public Policy

@@ Abstract
Introduction
In recent years, black-box optimization (BBO) has emerged as a critical approach in addressing complex decision-making challenges across various domains, particularly when dealing with objective functions that are difficult to analyze or explicitly define. Unlike traditional optimization methods that require gradient information or explicit mathematical formulations, black-box optimization treats the objective function as a “black box” that can be queried for function values but offers no additional information about its structure (Pardalos et al. 2021). This methodology is especially valuable for designing public policy, such as police redistricting (Zhu et al. 2020; Zhu, Wang, and Xie 2022), site selection for emergency service systems (Xing and Hua 2022), hazard assessment (Xie et al. 2021) and public healthcare policymaking (Chandak et al. 2020). Policymakers and researchers in these areas frequently encounter optimization problems embedded within complex human systems, where decision evaluations are inherently implicit, and conducting them can be resource-intensive. Black-box optimization, therefore, stands out as a potent tool for navigating through these intricate decision spaces, providing a means to optimize outcomes without necessitating a detailed understanding of the underlying objective function’s analytical properties.
@@ Introduction
However, the broader application prospects of BBO in public policymaking are hampered due to two major hurdles: (1) Defining the feasible region or setting clear constraints for decisions is inherently complicated for real human systems. Policymakers often encounter a myriad of both explicit and implicit rules when making an optimal decision, adding a significant layer of complexity to the optimization process. For instance, in police districting, police departments often organize their patrol forces by dividing the geographical region of a city into multiple patrol areas called zones. Their goal is to search for the optimal districting plan that minimizes the workload variance across zones (Larson 1974; Larson and Odoni 1981; Zhu et al. 2020; Zhu, Wang, and Xie 2022). As shown in Figure 1, a well-conceived plan necessitates each zone to adhere to certain shape constraints (e.g., contiguity and compactness) that are analytically challenging to formulate (Shirabe 2009), while also taking socio-economic or political considerations into account (e.g., ensuring fair access to public facilities). This creates a web of implicit constraints (Choi et al. 2000; Audet, Caporossi, and Jacquet 2020) that are elusive to define clearly and have a complex high-dimensional structure (e.g., manifold shape), making the assessment of feasible region nearly as expensive as evaluating the objective itself (Gardner et al. 2014). (2) The decisions for public policy are usually high-dimensional, presenting a significant computational hurdle to utilizing traditional BBO methods (Luong et al. 2019; Wan et al. 2021; Binois and Wycoff 2022). For example, the redistricting problem can be formulated by mixed-integer programming, grappling with hundreds or even thousands of decision variables even for medium-sized service systems (Zhu et al. 2020; Zhu, Wang, and Xie 2022).
@@
Despite the difficulty in formulating constraints for decisions in public policymaking, the rich repository of historical decisions adopted by the practitioners, combined with the increasingly easier access to human systems (van Veenstra and Kotterink 2017; Yu et al. 2021), offer a wealth of decision samples. Collecting these samples might involve seeking guidance from official public entities on decision feasibility or generating decisions grounded in domain expertise, bypassing the need to understand the explicit form of the constraints. These readily available decisions harbor implicit knowledge that adeptly captures the dynamics of implicit constraints, providing a unique opportunity to skillfully address these issues. This inspires us to develop a conditional generative representation model that maps the feasible region in the original space to a lower-dimensional latent space, which encapsulates the key pattern of these implicit constraints. As a result, the majority of existing BBO methods can be directly applied to solve the original optimization problem in this latent space without constraints.
@@
In this paper, we aim to solve Implicit-Constrained Black-Box Optimization (ICBBO) problems. In these problems, the constraints are not analytically defined, however the feasibility of a given decision can be easily verified. We introduce a new approach called Conditional And Generative Black-box Optimization (CageBO). Using a set of labeled decisions as feasible or infeasible, we first construct a conditional generative representation model based on the conditional variational autoencoder (CVAE) (Sohn, Lee, and Yan 2015) to explore and generate new feasible decisions within the intricate feasible region. To mitigate the impact of potentially poorly generated decisions, we adopt a post-decoding process that aligns these decisions closer to the nearest feasible ones. Furthermore, we incorporate this conditional generative representation model as an additional surrogate into our black-box optimization algorithm, allowing for a two-way mapping between the feasible region in the original space and an unconstrained latent space. As a result, while objective function is evaluated in the original space, the black box optimization algorithm is performed in the latent space without constraints. We prove that our proposed algorithm can achieve a no-regret upper bound with a judiciously chosen number of observations in the original space. Finally, we validate our approach with numerical experiments on both synthetic and real-world datasets, including applying our method to address large-scale redistricting challenges within police operation systems in Atlanta, Georgia, demonstrating its significant empirical performance and efficiency over existing methods.
@@
    We formulate a novel class of optimization problems in public policy, termed Implicit Constrained Black-Box Optimization (ICBBO), and develop the CageBO algorithm that effectively tackles high-dimensional ICBBO problems.
@@
    We introduce a conditional generative representation model that constructs a lower-dimensional, constraint-free latent space, enabling BBO algorithms to efficiently explore and generate feasible candidate solutions. This model supports the development of innovative public policy decisions by facilitating the exploration of new and effective policy options.
@@
    By establishing a no-regret expected cumulative regret bound, we demonstrate that our CageBO algorithm can consistently identify the global optimal solution, ensuring the development of near-optimal public policy decisions.
@@
    We apply our CageBO algorithm to the Atlanta police redistricting problem using both synthetic and real datasets. Our results demonstrate its superior performance compared to baseline methods, highlighting its potential to significantly enhance efficiency in real-world public policy applications.
@@
Related work. Black-box optimization, a.k.a. zeroth-order optimization or derivative-free optimization, is a long-standing challenging problem in optimization and machine learning. Existing work either assumes the underlying objective function is drawn from some Gaussian process (Williams and Rasmussen 2006) or some parametric function class (Dai et al. 2022; Liu and Wang 2023). The former one is usually known as Bayesian optimization (BO), with the Gaussian process serving as the predominant surrogate model. BO has been widely used in many applications, including but not limited to neural network hyperparameter tuning (Kandasamy et al. 2020; Turner et al. 2021), material design (Ueno et al. 2016; Zhang, Apley, and Chen 2020), chemical reactions (Guo, Ranković, and Schwaller 2023), and public policy (Xing and Hua 2022).
@@
In numerous real-world problems, optimization is subject to various types of constraints. Eriksson proposes scalable BO with known constraints in high dimensions (Eriksson and Poloczek 2021). Letham explores BO in experiments featuring noisy constraints (Letham et al. 2019). Gelbart pioneered the concept of BO with unknown constraints (Gelbart, Snoek, and Adams 2014), later enhanced by Aria through the ADMM framework (Ariafar et al. 2019). Their constraints are unknown due to uncertainty but can be evaluated using probabilistic models. In addition, Choi and Audet study the unrelaxable hidden constraints in a similar way (Choi et al. 2000; Audet, Caporossi, and Jacquet 2020), where the feasibility of a decision can be evaluated by another black-box function. In contrast, the implicit constraints are often unknown due to the lack of analytical formulations in many public policy-making problems.
@@
Building on latent space methodologies, Varol presented a constrained latent variable model integrating prior knowledge (Varol et al. 2012). Eissman (Eissman et al. 2018) presents a VAE-guided Bayesian optimization algorithm with attribute adjustment. Deshwal and Doppa focus on combining latent space and structured kernels over combinatorial spaces (Deshwal and Doppa 2021). Maus further investigates structured inputs in local spaces (Maus et al. 2022), and Antonova introduces dynamic compression within variational contexts (Antonova et al. 2020). However, it’s worth noting that none of these studies consider any types of constraints in their methodologies.
In public policy-making, implicit constraints are often a critical consideration in queueing service models. The pioneering work on the hypercube queueing model, introduced in Larson’s paper (Larson 1974) and further developed in his book (Larson and Odoni 1981), established a foundational framework for analyzing spatial queueing systems, particularly in the context of emergency services like police and ambulance operations. This model is frequently employed as a black-box performance measure for public resource deployment (de Souza et al. 2015; Ansari, McLay, and Mayorga 2017; Zhu et al. 2020). However, real-world service systems are often subject to various constraints on feasible policies. These constraints, such as fairness (Argyris, Karsu, and Yavuz 2022) and continuity (Zhu, Wang, and Xie 2022), are typically implicit and lack explicit analytical formulations, posing challenges for effective optimization.
Preliminaries
@@
Bayesian optimization. The BO algorithms prove especially valuable in scenarios where the evaluation of the objective function is costly or time-consuming, or when the gradient is unavailable. This approach revolves around constructing a surrogate model of the objective function and subsequently employing an acquisition function based on this surrogate model to determine the next solution for evaluation. For the minimization problem, a popular choice of surrogate model is the Gaussian process with the lower confidence bound (LCB) (Srinivas et al. 2009) serving as the acquisition function.
@@
Conditional Generative Representation
@@
Suppose there exists a joint distribution between the decision variable $x\in\mathcal{X}$, which can be defined in a high-dimensional hybrid space of discrete and continuous variables, and a continuous latent variable $z\in\mathcal{Z}\subseteq\mathbb{R}^{d}$, conditioned on the feasibility criterion $c\in\{0,1\}$. We model the conditional distributions $q_{\phi}(z|x,c)$, $p_{\theta}(x|z,c)$, and the conditional prior $p(z|c)$ through the use of neural networks as outlined in (Pinheiro Cinelli et al. 2021). Here, $\phi$ and $\theta$ represent the neural network weights associated with each respective distribution. Since it is intractable to directly optimize the marginal maximum likelihood of $x$, the evidence lower bound (ELBO) (Jordan et al. 1999) of the log-likelihood is derived as follows
$\mathcal{L}_{\mathrm{ELBO}}$ $=w(c)\underset{q_{\phi}(z|x,c)}{\mathbb{E}}\left[\log p_{\theta}(x|z,c)\right]$
$\quad-\eta\mathrm{D}_{\mathrm{KL}}\left(q_{\phi}(z|x,c)\|p(z|c)\right),$ (1)
@@
The post-decoding process is initiated only for decoder-generated decisions $x$ that are infeasible. Feasible decisions are directly incorporated into the observed feasible set $\mathcal{D}_{f}$. Through iteratively expanding this observed feasible set, we enhance the accuracy of this process by improving the coverage of $\mathcal{D}_{f}$ on the underlying feasible region $\mathcal{S}$. This is further detailed in the ablation study in the appendix.
Surrogate Model
@@
We use a GP as our surrogate model of the indirect objective function $g$, denoted by $\text{GP}({\mu},{k};\mathcal{Z})$. In our problem, the mean function can be written as ${\mu(z)=\mathbb{E}[g(z)]}$. In addition, we adopt the Matérn kernel (Seeger 2004) as the kernel function $k(z,z^{\prime})=\mathbb{E}[(g(z)-\mu(z))(g(z^{\prime})-\mu(z^{\prime}))]$, which is widely-used in BO literature. The main advantage of the GP as a surrogate model is that it can produce estimates of the mean evaluation and variance of a new latent variable, which can be used to model uncertainty and confidence levels for the acquisition function described in the following. Note that the latent variable $z$ is assumed to follow a Gaussian prior in the latent decision model, which aligns with the assumption of the GP model that the observed latent variables $Z$ follow the multivariate Gaussian distribution.
@@
Figure 3: Solution paths of the same optimization problem suggested by our CageBO algorithm in the (a) original and (b) latent spaces, respectively. The objective in this illustrative example is to minimize the Ackley function subject to an ”implicit” constraint (a circle) in a 2-dimensional decision space.
@@
We assume the distance between any two points in $\mathcal{Z}$ can be upper bounded by their distance in $\mathcal{X}$, i.e., $\forall x,x^{\prime}\in\mathcal{X},\|\mathtt{enc}(x)-\mathtt{enc}(x^{\prime})\|_{2}\leq C_{p}\|x-x^{\prime}\|_{2}$. We further assume that function $g:\mathcal{Z}\mapsto\mathbbm{R}$ is drawn from a Gaussian Process prior and it is $C_{g}$-Lipschitz continuous, i.e., $\forall z,z^{\prime}\in\mathcal{Z},|g(z)-g(z^{\prime})|\leq C_{g}\|z-z^{\prime}\|_{2}$. This assumption is the standard Gaussian process assumption for Bayesian optimization (Srinivas et al. 2010). Here $C_{p}$ and $C_{g}$ are universal constants. We use big $\tilde{O}$ notation to omit any constant and logarithmic terms. Now we are ready to state our main theoretical result.
@@
We present an upper bound of expected cumulative regret of our Algorithm 1. This is a no-regret algorithm since $\lim_{T\rightarrow\infty}\mathbb{E}[R_{T}]/T=0$ which means our algorithm is able to find the global optimal solution by expectation. The bound has two terms. The first term follows from GP-UCB (Srinivas et al. 2010) where the maximum information gain depends on the choice of kernel used in the algorithm. For linear kernel, $\gamma_{T}=O(d\log T)$ and for squared exponential kernel, $\gamma_{T}=O(\log T)^{d+1}$. The second term is the regret term incurred by the post decoder $\psi_{\mathcal{D}_{f}}$ when needed, which is also sublinear in $T$ if $n$ is chosen no larger than $T$. Due to page limit, full proof of Theorem 1 is shown in the appendix.
Experiments
Experimental Setup
We compare our CageBO algorithm against three baseline approaches that can be used to address ICBBO problems. These baselines include (1) vanilla version of Bayesian optimization (BO) (Jones, Schonlau, and Welch 1998), (2) simulated annealing (SA) (Kirkpatrick, Gelatt Jr, and Vecchi 1983), (3) approximated Mixed Integer Linear Programming (MILP) (Zhu, Wang, and Xie 2022), (4) a VAE-guided Bayesian optimization (VAE-BO) (Eissman et al. 2018), and (5) a variant of constrained VAE-guided Bayesian optimization (CON-BO) (Griffiths and Hernández-Lobato 2020). We train the latent decision model in our framework with the Adam optimizer (Kingma and Ba 2015). Both CageBO, VAE-BO, and CON-BO are trained and performed under an identical environment. MILP is only compared in the redistricting problems because their implicit constraints can be approximated by a set of linear constraints, albeit with the trade-offs of adding auxiliary variables and incurring computational expenses. Each method is executed 10 times across all experiments to determine the $95\%$ confidence interval of their results.
@@
Synthetic Results
(a) Keane’s bump function
(b) Michalewicz function
@@
We consider minimizing (a) the $30$-d Keane’s bump function (Keane 1994) and (b) the $30$-d Michalewicz function (Molga and Smutnicki 2005), both of which are common test functions for constrained optimization. To obtain samples from implicit constraints, the conundrum we aim to address via our methodology, we first generate $n=2,000$ samples from the standard uniform distribution in the $10$-dimensional latent space, then decode half of them through a randomly initialized decoder and mark those as feasible. We define the feasibility oracle to return 1 only if the input solution is matched with a feasible solution. We scale the samples accordingly so that the test functions can be evaluated under their standard domains, i.e., the Keane’s bump function on $[0,10]^{d}$ and the Michalewicz function on $[0,\pi]^{d}$.
@@
Case Study: Police Redistricting
@@
In this equation, ${\tau}_{j}$ represents the average travel time within district $j$, which can be estimated using the hypercube queuing model (Larson 1974), which can be regarded as a costly black-box function. Assuming $\mu$ represents the uniform service rate across all districts and $\lambda_{j}$ denotes the arrival rate in district $j$. Essentially, $\rho_{j}$ quantifies the cumulative working duration per unit time for police units in district $j$. For clarity, a value of $\rho_{j}=10$ implies that the combined working time of all police units in district $j$ counts to 10 hours or minutes in every given hour or minute.
(a) MILP
(b) SA
(c) VAE-BO
(d) CageBO
(e) MILP
(f) SA
(g) VAE-BO
(h) CageBO
Figure 5: Districting plans for the police redistricting problem in a synthetic $6\times 6$ grid and a real-world scenario in Atlanta, Georgia. The workload in each region is indicated by the depth of the color. The CageBO plan achieves the lowest workload variance among all districts, indicating an optimal distribution of workload.
@@
Atlanta police redistricting. For the police zone design in Atlanta, there are 78 police service regions that need to be divided into 6 zones, with the decision variable represented as $x\in\{0,1\}^{78\times 6}$. The redistricting of Atlanta’s police zones is constrained by several potential implicit factors. These include contiguity and compactness, as well as other practical constraints that cannot be explicitly defined. One such implicit constraint is the need for changes to the existing plan should be taken in certain local areas. A drastic design change is undesirable because: (1) A large-scale operational change will result in high implementation costs. (2) A radical design change will usually face significant uncertainties and unpredictable risks in future operations. The arrival rate $\lambda_{l}$ and the service rate $\mu$ are estimated using historical 911-calls-for-service data collected in the years 2021 and 2022 (Zhu and Xie 2019, 2022; Zhu, Wang, and Xie 2022). We generate an initial data set $\mathcal{D}$ by simulating $n=10,000$ labeled decisions which are the neighbors of the existing plan with changes in certain local areas.
@@
Figure 4(d) displays the convergence of our algorithm in comparison to other baseline methods, showing a consistent reduction in workload variance. Notably, the redistricting plan created by our CageBO algorithm most closely aligns with the actual heuristic policy applied by policymakers, wherein the zone with the highest arrival rate is assigned the largest workload, and the workloads of other zones are well-balanced. Further discussion on fair districting can be found in the appendix. As shown in Figure 5, the plan produced by our CageBO algorithm achieves the lowest workload variance, surpassing the baseline algorithms. The superior performance of our CageBO algorithm highlights its capacity to offer critical managerial insights to policymakers, thereby facilitating informed decision-making in real-world applications.
Conclusion